If the general solution to a system of equations with infinite solutions is
(2t-2,5t+1,t), then what is the particular solution that goes with
t=0?

Found 2 solutions by ikleyn, Param:
Answer by ikleyn(52780)

.

Substitute t= 0 into the formulas in parentheses.



Answer by Param(1)

2t-2=x
5t+1=y
t=z
so, for these equations fill the value of t=0
2(0)-2=-2
5(0)+1=1
0=0

the value of x=-2
y=1
z=0